### v1

# Efficient Visual State Space Model for Image Deblurring

Abstract
Convolutional neural networks (CNNs) and Vision Transformers (ViTs) have achieved excellent performance in image restoration. ViTs typically yield superior results in image restoration compared to CNNs due to their ability to capture long-range dependencies and input-dependent characteristics. However, the computational complexity of Transformer-based models grows quadratically with the image resolution, limiting their practical appeal in high-resolution image restoration tasks. In this paper, we propose a simple yet effective visual state space model (EVSSM) for image deblurring, leveraging the benefits of state space models (SSMs) to visual data. In contrast to existing methods that employ several fixed-direction scanning for feature extraction, which significantly increases the computational cost, we develop an efficient visual scan block that applies various geometric transformations before each SSM-based module, capturing useful non-local information and maintaining high efficiency. Extensive experimental results show that the proposed EVSSM performs favorably against state-of-the-art image deblurring methods on benchmark datasets and real-captured images.
1 Introduction
Image deblurring aims to restore sharp images from blurry ones, attracting much attention due to the popularization of various cameras and hand-held imaging devices. This task is challenging as only the blurred images are available while lacking access to the blur and latent images.
Significant progress has been made due to the development of deep convolutional neural networks (CNNs) [40, 11, 5, 53, 3]. However, as the main operation in CNNs, the convolution is spatially invariant and spatially local. This does not capture the spatially variable properties of the image contents and cannot explore non-local information that is beneficial for deblurring.
In contrast to the convolution operation, the self-attention mechanism [44] in Transformers can capture global information by computing correlations between each token and all other tokens, which can extract better features for image deblurring. However, the self-attention mechanism (i.e., the scaled dot-product attention) entails quadratic space and time complexity regarding the number of tokens, which becomes unacceptable when handling high-resolution images. Although the local window-based methods [48, 26], transposed attention [54], and the frequency domain-based approximation [21] have been developed to reduce the computational cost, these approaches sacrifice their abilities to model non-local information [48, 26] and spatial information [54, 21], which thus affects the quality of restored images. Therefore, it is of great need to develop an efficient approach that can explore non-local information for high-quality deblurring performance while not significantly increasing the computational cost.
Recently, state space models (SSMs) [16, 15] have demonstrated significant potential in modeling long-range dependencies for natural language processing (NLP) tasks with linear or near-linear computational complexity. The improved SSM, specifically Mamba [13], develops a selective scan mechanism (S6) and can remember relevant information and ignore irrelevant contents while achieving linear computational complexity. This motivates us to utilize Mamba to explore useful non-local information for better image deblurring efficiently. However, as Mamba is designed for handling one-dimensional (1D) sequences, it requires first flattening the image data to a 1D image sequence if straightforwardly applying Mamba to visual tasks [12, 7]. This disrupts the spatial structure of the image, making it difficult to capture local information from various adjacent pixels. Several approaches adopt a multi-direction scan mechanism to utilize the state space model in visual applications [27, 17, 37]. However, the multi-direction scan mechanism significantly increases the computational cost.
In this paper, we propose an effective and efficient visual state space model for image deblurring. We find that existing visual state space models mostly adopt several fixed-direction scanning for feature extraction, which may not explore non-local information adaptively and lead to higher computational costs, We thus develop a simple yet effective scanning strategy that captures non-local spatial information while maintaining low computational costs. Specifically, we only scan the input feature in one direction but employ a simple geometric transformation before each scan, which effectively and adaptively explores useful information with a minimal increase in computational costs.
The main contributions are summarized as follows. First, we propose a simple yet effective visual state space model that efficiently restores high-quality images. Second, we develop an effective and efficient scanning strategy that captures non-local spatial information while maintaining low computational costs. Finally, we quantitatively and qualitatively evaluate the proposed method on benchmark datasets and real-world images and show that it performs effectively and efficiently against state-of-the-art methods.
2 Related Work
Deep CNN-based image deblurring methods.
In recent years, significant progress has been made in image deblurring through deep CNN-based methods [31, 40, 11, 55, 53, 5, 3]. In [31], a deep CNN is proposed based on a multi-scale framework that directly estimates clear images from blurry ones. Tao et al. [40] introduce an effective scale recurrent network to enhance the utilization of information from each scale within the multi-scale framework. Furthermore, Gao et al. [11] propose a selective network parameter-sharing method to enhance further the methods proposed in [31, 40]. Additionally, generative adversarial networks (GANs) have been widely applied in image deblurring [22, 23], aiming to enhance the quality of deblurred results by generating realistic and sharp images.
Due to the limited improvement in performance with the utilization of additional scales, Zhang et al. [55] propose an effective network that adopts a multi-patch strategy for image deblurring. The deblurring process is executed step-by-step, enabling the network to refine the output progressively. To further exploit the features extracted at different stages, Zamir et al. [53] introduce a cross-stage feature fusion technique, aiming to enhance the overall performance of the deblurring method. To mitigate the problem of high computational cost associated with multi-scale frameworks, Cho et al. [5] present a multi-input and multi-output network architecture, reducing the computational burden while maintaining the deblurring performance. Chen et al. [3] analyze the baseline modules and propose simplified versions to improve the efficiency of image restoration. However, as the convolution operation is spatially invariant and spatially local, it cannot effectively model the global and spatially-variant information, limiting its ability to achieve better image restoration.
Transformer-based image deblurring methods.
As Transformers can establish long-range dependencies and effectively model global information, it has achieved significant advancements in various high-level vision tasks such as image classification [28, 47], object detection [1, 56], and semantic segmentation [57, 51]. Researchers have extended its application to image super-resolution [26], image deblurring [54, 21], and image denoising [2, 48]. The Transformer’s self-attention mechanism requires quadratic computational complexity, which is unacceptable for image restoration tasks with high-resolution images. To reduce the computational complexity of Transformers, Zamir et al. [54] propose an efficient Transformer model that computes scaled dot-product attention in the feature depth domain. Tsai et al. [42] simplify the self-attention calculation by constructing intra- and inter-strip tokens to replace global attention. Wang et al. [48] introduce a Transformer based on a UNet architecture that applies non-overlapping window-based self-attention for single image deblurring. Kong et al.  [21] propose a frequency domain-based Transformer and achieve state-of-the-art results. Although these approaches employ various strategies to reduce computational complexity, they cannot characterize long-range dependencies and non-local information effectively. In contrast, we develop an efficient visual state space model that can explore useful non-local information with low computational costs.
State space models.
State Space Models (SSMs) have been a time series analysis and modeling cornerstone for decades. Recent methods [16, 15] have adopted SSMs to capture long-range dependencies for sequence modeling. SSM-based methods can be computed efficiently using recurrence or convolution, with linear or near-linear computational complexity. Gu et al. [14] propose a framework (HiPPO) that poses the abstraction of optimal function approximation concerning time-varying measures. The method [15] develops a linear state-space layer to handle long-range dependency problems. To address the issue of high computational and memory requirements induced by state representation, S4 [16] proposes a method of normalizing parameters into a diagonal structure. Furthermore, Mamba [13] introduces a selective scanning layer with dynamic weights, showcasing significant potential in natural language processing. To apply SSMs for visual tasks, recent methods [27, 37, 17] adopt multi-direction scanning strategies, which will increase the computational cost. In contrast, we propose an efficient visual scan block that employs a geometric transformation before each scan to achieve non-local information exploration with high efficiency.
Figure 1: Efficient visual state space model. To efficiently restore high-quality images with SSMs, we propose an effective EVSS module that involves an efficient EVS block and an efficient EDFFN block. A geometric transformation is employed at the beginning of each EVS block to facilitate more useful information exploration in the following selective scan with a minimal increase in computational costs.
3 Efficient Visual State Space Model
Our goal is to present an effective and efficient method to explore the properties of state space models for high-quality image deblurring. To this end, we propose an efficient visual state space model (EVSSM) that can explore more non-local information for visual tasks with a minimal increase in computational costs.
3.1 Overall architecture
As shown in Figure 1, the overall architecture of the proposed efficient visual state space model is based on a hierarchical encoder-decoder framework [21]. Given a blurry image $I_{blur}\in\mathbb{R}^{H\times W\times 3}$, we first employ a $3\times 3$ convolution layer to obtain the shallow feature $F_{s}\in\mathbb{R}^{H\times W\times C}$, where $H\times W$ denotes the spatial dimension, and $C$ is the number of feature channels. Then, the shallow feature $F_{s}$ is put into a 3-level symmetric encoder-decoder network. The encoder/decoder at each level is composed of several efficient vision state space (EVSS) modules (see Section 3.2). For the encoder/decoder at level-$l$, the input feature is progressively processed by each EVSS module and the intermediate feature ${F^{l}_{enc}}$ / ${F^{l}_{dec}}$ $\in\mathbb{R}^{\frac{H}{2^{l-1}}\times\frac{W}{2^{l-1}}\times{2^{l-1}}C}$ ($l=1,2,3$ in this work) is generated. We then employ the bilinear interpolation and $1\times 1$ convolution to achieve upsampling and downsampling and add the skip connection between the encoder and decoder at each level. Finally, a $3\times 3$ convolution layer is applied to the feature ${F^{3}_{dec}}$ to generate the residual image $R\in\mathbb{R}^{H\times W\times 3}$.
The restored image $I_{deblur}$ is obtained by $I_{deblur}=R+I_{blur}=\mathcal{N}(I_{blur})+I_{blur}$, and $\mathcal{N}$ denotes the proposed encoder-decoder network regularized by minimizing the following loss function:
$\mathcal{L}=\lVert I_{deblur}-I_{gt}\rVert_{1}+\lambda\lVert\mathcal{F}(I_{deblur})-\mathcal{F}(I_{gt})\rVert_{1},\$ (1)
where $\mathcal{F}$ denotes the discrete Fourier transform and the weight parameter $\lambda$ is empirically set as 0.1.
3.2 Efficient vision state space module
State space model.
A state space model is a mathematical framework commonly used in time series analysis and control systems. The state equation describes the evolution of an underlying system over time, representing the relationship between the system’s hidden states and their temporal dynamics. The input signal $x(t)$ is mapped to the output response $y(t)$ through the hidden state $h(t)$. It is typically modeled as a set of first-order difference or differential equations:
$h^{{}^{\prime}}(t)=Ah(t)+Bx(t),~{}~{}~{}~{}y(t)=Ch(t)+Dx(t),$ (2)
where $A,B,C$, and $D$ are learnable weight matrices.
To this end, the state equation can be discretized using the zero-order hold (ZOH) technique:
$\begin{split}&h_{t}=\bar{A}h_{t-1}+\bar{B}x_{t},~{}~{}~{}~{}y_{t}=Ch_{t}+Dx_{t},\\ &\bar{A}=e^{\Delta A},~{}~{}~{}~{}\bar{B}=(\Delta A)^{-1}(e^{\Delta A}-I)\cdot\Delta B.\end{split}$ (3)
Based on (3), Mamba [13] proposes a mechanism of selective scanning (S6) to achieve input-dependent weights and linear computational complexity simultaneously. Using a state space model for NLP tasks poses no issue since natural language data is inherently a causal sequence. However, visual tasks present significant challenges because visual data is fundamentally non-sequential and contains spatial information such as local textures and global structures. As S6 is a recursive process, when processing an input at the current timestep $t$, it can only utilize information from previous timesteps and not consider information from future timesteps.
Efficient visual scan (EVS) block.
A straightforward approach to solve this problem is to scan the visual data in different directions (e.g., forward and backward). However, this strategy increases the computational cost by a significant factor. For example, the computational cost of the scanning in VMamba [27] is 4×4$\times$4 × higher than that of Mamba [13] due to its strategy of performing bidirectional scanning in the longitudinal and transverse directions. Then, a natural question is whether it can reasonably process visual data using a state space model with a minimal increase in computational costs.
The answer is YES. We develop an efficient vision state space model that explores more useful information with a minimal increase in computational costs. The key is the proposed EVS block. Instead of scanning in multiple directions, we scan only in one direction and apply one geometric transformation (e.g., flip, and transposition) to the input before each scan. Due to the translation-invariant property of convolution, the geometric transformation does not affect the convolution itself but only influences the process of selective scanning. Specifically, for each EVS block, assuming that it is in the $i$-th EVSS module of the whole network, we first transpose or flip the input feature $F_{in}$ as:
$\begin{split}&G=\begin{cases}Transpose(F_{in})&\text{if }i~{}\%~{}2=0,\\ Flip(F_{in})&\text{if }i~{}\%~{}2=1.\\ \end{cases}\end{split}$ (4)
Here, we flip along both the horizontal and vertical axes of the feature in $Flip(\cdot)$. According to (4), the image feature will be automatically restored to the original spatial structure after every 4 EVSS modules. In particular, if the total number of EVSS modules in our network is not divisible by 4, we can restore the original spatial structure by applying the corresponding inverse transformations, as both flip and transposition are reversible. In this way, our EVSS module effectively solves the above question, avoiding any additional computational burden other than efficient geometric transformations. Then, the selective scan can be formulated as follows:
$\begin{split}&X_{1},X_{2}=\mathrm{split}(\mathrm{Conv}_{1\times 1}(G))\\ &\hat{X_{1}}=\mathrm{S6}(\mathrm{Reshape}(\sigma(\mathrm{Dconv_{3\times 3}}(X_{1}))))\\ &\hat{X_{2}}=\sigma(X_{2})\\ &F_{out}=\mathrm{Conv}_{1\times 1}(\mathrm{Reshape}(\mathcal{L}(\hat{X_{1}}))\cdot\hat{X_{2}}),\end{split}$ (5)
where $\mathrm{Conv}_{1\times 1}(\cdot)$ denotes a convolutional layer with the filter size of $1\times 1$ pixel, $\mathrm{DConv}_{3\times 3}(\cdot)$ denotes a depth-wise convolutional layer with the filter size of $3\times 3$ pixels, $\mathcal{L}(\cdot)$ denotes a normalization layer, $\mathrm{split}(\cdot)$ splits the image features in the channel dimension, $\sigma$ denotes the GeLU activation, and S6 denotes the selective scanning mechanism proposed by [13]. Figure 1(b) shows the detailed network architecture.
Efficient discriminative frequency domain-based FFN (EDFFN).
To effectively and efficiently transform the features from the EVSS module, we develop an efficient discriminative frequency domain-based FFN. The FFN part is typically the core component of deep learning models which can help the latent clear image reconstruction [54, 21]. FFTformer [21] develops a discriminative frequency domain-based FFN that adaptively determines which frequency information should be preserved. However, this increases the computational cost when performing the frequency domain-based operations. In contrast to DFFN which applies frequency-domain based operations in the middle of the FFN network, our approach is to perform frequency-domain screening on the features at the final stage of the FFN network. Figure 1(c) shows the detailed network architecture.
4 Experimental Results
4.1 Datasets and implementation
Datasets.
Following existing state-of-the-art methods, we evaluate our approach on the commonly used GoPro dataset by Nah et al. [31], HIDE dataset by Shen et al. [36], and RealBlur dataset by Rim et al. [35]. The GoPro dataset contains $2103$ images for training and $1111$ images for testing. The HIDE dataset includes $2025$ images mainly about humans for testing. The RealBlur dataset [35] contains RealBlur-J and RealBlur-R subsets generated by different post-processing strategies. It uses $182$ scenes for training and $50$ scenes for testing. For fair comparisons, we follow the protocols of these datasets to evaluate our method.
Implementation details.
The shallow feature $F_{s}$ has a channel number of 48 and the numbers of EVSS modules in the encoder/decoder from level-$1$ to level-$3$ are [6, 6, 12]. We use the ADAM optimizer [20] with default parameters in the training process. We use the data augmentation method with the flipping and rotation operations to generate training data. We apply the progressive training similar to but simpler than [54]: the training starts with the patch size of $128\times 128$ pixels and batch size $64$ for $300,000$ iterations, where the learning rate gradually reduces from $1\times 10^{-3}$ to $1\times 10^{-7}$. Then the patch size is enlarged to $256\times 256$ pixels with $16$ batches for $300,000$ iterations where the learning rate is initialized as $5\times 10^{-4}$ and decreases to $1\times 10^{-7}$. The learning rate is updated based on the Cosine Annealing scheme. Unless otherwise specified, all experiments are conducted with the PyTorch framework on NVIDIA RTX 4090 GPUs. The training code and test model are available at https://github.com/kkkls/EVSSM.
Table 1: Quantitative evaluations of the proposed method against state-of-the-art ones on the GoPro dataset [31] in terms of PSNR and SSIM.
Method SRN [40] DMPHN [55] MIMO-UNet+ [5] MPRNet [53] MAXIM [43] Uformer [48] Restormer [54]
PSNR (dB) 30.26 31.20 32.45 32.66 32.86 33.06 32.92
SSIM 0.9342 0.9453 0.9567 0.9589 0.9616 0.9670 0.9611
Method Stripformer [42] Restormer-local [6] NAFNet [3] GRL [25] FFTformer [21] CU-mamba [7] EVSSM (ours)
PSNR (dB) 33.08 33.57 33.71 33.93 34.21 33.53 34.50
SSIM 0.9624 0.9656 0.9668 0.9680 0.9692 0.9650 0.9712
Table 2: Quantitative evaluations of the proposed method against state-of-the-art ones on the real-world dataset [35] in terms of PSNR and SSIM.
Dataset Method DeblurGAN-v2 [23] SRN [40] MIMO-Unet+ [5] BANet [41] DeepRFT+ [30] Stripformer [42] FFTformer [21] EVSSM (ours)
RealBlur-J PSNR (dB) 29.69 31.38 31.92 32.00 32.19 32.48 32.62 34.15
SSIM 0.8703 0.9091 0.9190 0.9230 0.9305 0.9290 0.9326 0.9450
RealBlur-R PSNR (dB) 36.44 38.65 - 39.55 39.84 39.94 40.11 41.04
SSIM 0.9347 0.9652 - 0.9710 0.9721 0.9739 0.9732 0.9770
Table 3: Quantitative evaluations of the proposed method against state-of-the-art ones on the HIDE dataset [36] in terms of PSNR and SSIM.
Method SRN [40] DMPHN [55] SAPHN [39] MIMO-UNet+ [5] MPRNet [53] Uformer [48] MPRNet-local [6]
PSNR (dB) 28.36 29.09 29.98 29.99 30.96 30.90 31.19
SSIM 0.9040 0.9240 0.9300 0.9304 0.9397 0.9530 0.9418
Method Restormer [54] Stripformer [42] Restormer-local [6] NAFNet [3] GRL [25] FFTformer [21] EVSSM (ours)
PSNR (dB) 31.22 31.03 31.49 31.31 31.65 31.62 31.97
SSIM 0.9423 0.9395 0.9447 0.9427 0.947 0.9455 0.9501
Refer to caption Refer to caption Refer to caption
(a) Blurred input (b) GT (c) MIMO-UNet+ [5] (d) Stripformer [42]
Refer to caption Refer to caption Refer to caption
(e) Restormer-local [6] (f) NAFNet [3] (g) FFTformer [21] (h) EVSSM (ours)
Figure 2: Deblurred results on the GoPro dataset [31]. Compared to the results in (c)-(g), the proposed method generates a better deburred image with clearer structures and detail in (h).
4.2 Comparisons with the state of the art
Evaluations on the GoPro dataset.
We first evaluate the performance of the proposed approach on the GoPro dataset [31]. We compare our method against state-of-the-art approaches, including CNN-based ones (SRN [40], DMPHN [55], MIMO-UNet+ [5], MPRNet [53], NAFNet [3]), Transformer-based ones (Uformer [48], Restormer [54], Stripformer [42], Restormer-local [6], GRL [25], and FFTformer [21]), MLP-based ones (MAXIM [43]), and SSM-based ones (CU-mamba [7]). We retrain or fine-tune the deep learning-based methods for fair comparisons if they are not trained on the benchmarks. As CU-mamba [7] does not provide training and test codes, for fair comparison, we compare our result with the reported result of [7]. We use the PSNR and SSIM as the evaluation metrics to measure the quality of each restored image. Table 1 shows the quantitative evaluation results, where our approach outperforms other methods with the highest PSNR and SSIM values.
Figure 2 shows evaluation results on the GoPro dataset [31]. Since CNN-based methods cannot effectively utilize global information, the images restored by [5, 3] still contain severe blur residuals as shown in Figure 2(c) and (f). Although Transformer-based methods can model the global context, the methods [42, 6, 21] have adopted various approximations to reduce the computational cost. This affects their ability to model the global information, and as a result, some main structures, e.g., tire, and wheel, are not restored well (see Figure 2(d), (e) and (g)). In contrast to existing Transformer-based methods, we propose a simple yet effective visual state space model for image deblurring, which is effective at exploring non-local information with low computational costs. As shown in Figure 2(h), our approach generates better results, where the text and tire are much clearer.
Evaluations on the RealBlur dataset.
Using the same protocols, we evaluate our approach on the real-world blurry dataset by Rim et al. [35]. Table 2 shows that the proposed method outperforms previous work significantly, improving the PSNR by at least $1.53$dB and $0.93$dB on the datasets of RealBlur-J and RealBlur-R, respectively. Figure 3 shows the visual comparisons on the RealBlur dataset, where our method generates the result with clearer characters and finer structural details.
Refer to caption Refer to caption Refer to caption
(a) Blurred input (b) GT (c) SRN [40] (d) DeblurGAN-v2 [23]
Refer to caption Refer to caption Refer to caption
(e) MIMO-UNet+ [5] (f) Stripformer [42] (g) FFTformer [21] (h) EVSSM (ours)
Figure 3: Deblurred results on the RealBlur dataset [35]. The results obtained by [40, 23, 5] in (c)-(e) are still blurry. For other results by [42, 21] in (f)-(g), fine-scale structures are not effectively recovered. In contrast, our method restores a better-deblurred image with clearer characters.
Table 4: Quantitative evaluations of the proposed method against state-of-the-art ones on the real-world dataset [46] for image deraining in terms of PSNR and SSIM.
Method DSC [29] MSPFN [19] RCDNet [45] DualGCN [10] SPDNet [52] Restormer [54] IDT [50] DRSformer [4] EVSSM (ours)
PSNR (dB) 34.95 43.43 43.36 44.18 43.20 47.98 47.35 48.54 48.78
SSIM 0.9416 0.9843 0.9831 0.9902 0.9871 0.9921 0.9930 0.9924 0.9951
Table 5: Quantitative evaluations of the proposed method against state-of-the-art ones on the RESIDE-6K dataset [24] for image dehazing in terms of PSNR and SSIM.
Method DCP [18] MSCNN [33] GFN [34] MSBDN [8] PFDN [9] FFA-Net [32] AECRNet [49] DehazeFormer [38] EVSSM (ours)
PSNR (dB) 17.88 20.31 23.52 28.56 28.15 29.96 28.52 31.45 31.66
SSIM 0.816 0.863 0.905 0.966 0.962 0.973 0.964 0.980 0.968
Evaluations on the HIDE dataset.
We examine our method on the HIDE dataset [36] in Table 3. Similar to state-of-the-art methods [21, 6], we directly use the models trained on the GoPro dataset for testing. The PSNR value of our approach is $0.35$dB higher than those from competing methods, showing that our method generalizes well, as the models are not trained on this dataset.
Evaluations on datasets for other image restoration tasks.
We also evaluate our method on the real-world dataset [46] for image deraining and RESIDE-6K dataset [24] for image dehazing. Tables 4-5 show that the proposed approach performs favorably against state-of-the-art methods on image deraining and dehazing datasets, which demonstrates the generalization ability of our method on other image restoration tasks.
5 Analysis and Discussion
We have shown that the proposed efficient visual state space model generates favorable results compared to state-of-the-art methods. In this section, we provide more analysis of the proposed method and discuss the effect of the main components. For the ablation studies in this section, we train our approach and all the baselines on the GoPro dataset using a batch size of $64$ and a patch size of $128\times 128$ pixels.
Effectiveness of the EVS block.
The core of our proposed EVS block is the geometric transformation (4) applied at the beginning of this block. To demonstrate the effect of the EVS block, we first remove the geometric transformation (one-direction for short, which performs scanning in one direction as in [13]) and train this baseline model using the same settings as ours. Table 6 shows that the PSNR of our approach is $0.14$dB higher than this baseline method. The state space model requires flattening the image feature into a one-dimensional sequence, which compromises the spatial structural information of the visual data. Compared to the baseline, our method with the geometric transformation can better explore non-local information. Meanwhile, the number of parameters and the FLOPs of our method are the same as those of the baseline, and their runtimes are also almost identical, which demonstrates the effectiveness of the proposed EVS block in improving the ability of the state space model for handling visual data with a minimal increase in computational costs.
Table 6: Effectiveness of the proposed EVS block, evaluated on the GoPro dataset [31].
Scanning mode      Parameters (M) FLOPs (G)      Runtime (ms) PSNR/SSIM
one-direction 17.1 126 87.9 33.99/0.9683
two-direction 18.3 135 122.7 34.08/0.9687
four-direction 20.1 148 182.6 34.05/0.9686
EVS (ours) 17.1 126 88.7 34.13/0.9690
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
(a) Blurred input (b) one-direction (c) two-direction (d) four-direction (e) EVS (ours)
Figure 4: Effect of the proposed EVS block. Our method with the proposed EVS block is more effective at recovering clearer structures.
In addition, we also compare with two baselines that respectively perform scanning in two directions (two-direction for short) and four directions [27] (four-direction for short). As Table 6 shows, although scanning in multi-directions can alleviate the limitation of the state space model on handling visual data, it leads to an increase in the number of network parameters and the computational complexity, resulting in significantly longer runtimes, cf. $122.7$ms for one-direction vs. $182.6$ms for four-direction vs. $88.7$ms for our approach. Note that the results of scanning in four-directions are slightly lower than those of scanning in two-directions. This is because the methods that scan in multiple directions simply employ the summation followed by normalization to fuse features extracted from different scanning directions. Thus, the multi-direction information is not effectively and fully utilized. In contrast to simultaneously scanning in multiple directions, our approach applies geometric transformations to the input feature at the beginning of each EVS block. This allows each scan to capture the contextual information from different directions and mitigates the increase in computational complexity and runtime. Figure 4 demonstrates that our approach can deblur images better than other methods, where the structures of the windows are restored well (see Figure 4(e)).
Table 7: Effectiveness of the geometric transformation, evaluated on the GoPro dataset [31].
Method Scanning in one direction Transpose Flip Parameters (M) FLOPs (G) Runtime (ms) PSNR/SSIM
w/o F&T 17.1 126 87.9 33.99/0.9683
w/o F 17.1 126 88.4 34.05/0.9686
w/o T 17.1 126 88.5 34.03/0.9684
EVSSM (ours) 17.1 126 88.7 34.13/0.9690
Table 8: Model complexity of the top-performance methods on the GoPro dataset, evaluated on images with the size of $256\times 256$ pixels. All the results are obtained on a machine with an NVIDIA RTX 3090 GPU.
Method DMPHN [55] IPT [2] MPRNet-local [6] Restormer-local [6] GRL [25] FFTformer [21] EVSSM (ours)
Parameters (M) 21.7 114 20.1 26.1 20.20 16.6 17.1
FLOPs (G) 234 376 760 155 1289 131 126
Runtime (ms) 143 298 95 286 518 132 89
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
(a) Blurred input (b) w/o F&T (c) w/o F (d) w/o T (e) EVSSM (ours)
Figure 5: Effect of the geometric transformations. Compared to the results in (b)-(d), our approach with both the filp and transpose transformations generates a clearer image in (e).
Effectiveness of the geometric transformation.
In the EVS block, we adopt two classical image geometric transformations in the proposed EVS block: flip and transpose. To demonstrate their effectiveness, we individually remove the flip transformation (w/o F for short), the transpose transformation (w/o T for short), and both flip and transpose transformations (w/o F&T for short). The comparison results in Table 7 demonstrate that applying the flip or transpose transformation can achieve better results, improving the PSNR by at least $0.04$dB. Our approach, which uses both flip and transpose transformations, outperforms all these baseline methods without significantly increasing the computational cost or runtime. The visual comparisons in Figure 5 further demonstrate the effectiveness of our approach, where the wheel recovered by our method is much clearer.
Model complexity.
We further examine the model complexity of the proposed approach and other top-performance methods in terms of model parameters, floating point operations (FLOPs), and average runtime. Table 8 shows that the proposed method has fewer FLOPs and is faster than the evaluated methods.
Limitations.
We develop an effective and efficient method that explores the the properties of the state space model for high-quality image restoration. However, we have only considered simple transformations such as flip and transpose so far. In future work, we will consider more powerful transformation methods, such as the polar coordinate transformation, to better characterize the spatial information of visual data with SSMs.
6 Conclusion
In this paper, we propose an efficient visual state space model for image deblurring. Specifically, we develop an efficient visual scan block, where we employ various geometric transformations before each scan to adapt SSMs to visual data. We show that compared to existing methods that scan along multiple directions simultaneously, our method is more effective at exploring non-local information without significantly increasing the computational cost. Extensive evaluations and comparisons with state-of-the-art methods demonstrate that our approach is much more efficient while achieving favorable performance.